OpenStudy (anonymous):

@Abhisar A stationary object is released from the point P with 3R from the centre of the moon of radius R and mass m.Find an expression in terms of G,m and R for the speed of the object when it hits the moon's surface.

OpenStudy (ookawaiioo):

easyyyyy

OpenStudy (anonymous):



OpenStudy (ookawaiioo):

give me a minute.....i will give you a formula to your answer

OpenStudy (anonymous):

ok...thanks...

OpenStudy (ookawaiioo):

\[K.E =Force * distance\] \[\frac{ 1 }{ 2 }*mass*v^{2} = G \frac{ Moon*mass }{ radius }\]

OpenStudy (ookawaiioo):

Solve for v (velocity)

OpenStudy (ookawaiioo):

That is your answer

OpenStudy (abhisar):

Gravitational Force = \(\huge \frac{Gmm_o}{r^2}\\=>acceleratio~due~to~gravity=\huge \frac{Gm}{r^2}\)

OpenStudy (abhisar):

Now substitute the value of r as 3r and find acceleration

OpenStudy (abhisar):

Then use \(V^2=U^2+2as\) to find V, Take u=0 s=3r-r=2r

OpenStudy (ookawaiioo):

damnnnn....... you smart.

OpenStudy (anonymous):

Thank you very much...

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)